### v1

# Hidden heavy flavor tetraquarks in the Born-Oppenheimer approximation

Abstract
The Born-Oppenheimer approximation is one of the very successful tools for solving the hydrogen atom problem. The experimental discovery of hidden heavy flavor tetraquarks, $Q\bar{Q}q\bar{q}$ ($Q=c,b$ and $q=u,d,s$), provides great possibilities for the hydrogen-bond-like structure of the Quantum Chromodynamics version. In this work, considering that the colors of $Q\bar{Q}$ and $q\bar{q}$ are both 8, the color-spin wave functions of the tetraquark formed by the color coupling $8\otimes 8\rightarrow 1$ are constructed. Then we use the Born-Oppenheimer approximation to calculate the mass spectra of hidden heavy flavor tetraquarks and evaluate their spatial sizes. The results show that some of the hidden heavy flavor exotic hadrons discovered experimentally can be well explained as this type of hydrogen bond tetraquark structure. In addition, some candidates for tetraquark bound states formed by color-spin interactions are predicted and may be compact tetraquark states.
preprint: APS/123-QED
I Introduction
Quantum chromodynamics (QCD) provides a rich framework for understanding hadrons, extending beyond the conventional mesons (one quark, one antiquark) and baryons (three quarks) to encompass exotic configurations such as multiquark states, hybrids, and glueballs Godfrey and Isgur (1985); Klempt and Zaitsev (2007); Lebed et al. (2017). Among the multiquark states, the study of tetraquark states (composed of two quarks and two antiquarks) have received significant attention Jaffe (1977a, b). Notably, certain hidden heavy flavor tetraquarks, $Q\bar{Q}q\bar{q}$ ($Q=c,b$ and $q=u,d,s$), can be interpreted as analogues of hydrogen bonds within QCD version. This allows the interaction between heavy quark-antiquark pairs and light quarks can be treated using the Born-Oppenheimer (BO) approximation Esposito et al. (2013); Braaten et al. (2014); Maiani et al. (2019a, b); Mutuk (2024); Berwein et al. (2024, 2024); Germani et al. (2025). This approach decouples the heavy quark dynamics from the light quark dynamics, enabling systematic studies of tetraquark spectra and structure, based on the fact that the heavy quarks are much heavier than the light quarks.
In 2003, the Belle Collaboration reported the famous narrow charmonium-like state $X(3872)$ in the $\pi^{+}\pi^{-}J/\psi$ mass spectrum of the exclusive decay process $B^{\pm}\rightarrow K^{\pm}\pi^{+}\pi^{-}J/\psi$ Choi et al. (2003). The discovery of the $X(3872)$ confirmed the existence of exotic hadron states and opened up a new direction for hadron physics research. On the theoretical side, there are various explanations for the nature of the $X(3872)$, such as the tetraquark state Maiani et al. (2005); Matheus et al. (2007); Chen et al. (2016); Wang (2024), molecular state Wong (2004); AlFiky et al. (2006); Liu et al. (2008); Guo et al. (2013), $P$-wave charmonium Barnes and Godfrey (2004); Bali et al. (2011); Mohler et al. (2013); Aliev and Savcı (2015), and threshold effects Bugg (2004); Hanhart et al. (2007); Guo et al. (2015). Subsequently, more hidden heavy flavor exotic hadrons were discovered, many of which were considered to be candidates for tetraquark states and attracted much attention from particle physicists, such as $X(4140)$ Aaltonen et al. (2009); Padmanath et al. (2015); Luo et al. (2017), $Y(3940)$ Abe et al. (2005); Ebert et al. (2006); Lebed and Polosa (2016), $Z_{b}(10610)$ and $Z_{b}(10650)$ Bondar et al. (2012); Guo et al. (2011); Leskovec et al. (2019); see reviews for details Karliner et al. (2018); Guo et al. (2018); Brambilla et al. (2020); Chen et al. (2023). So far, although the understanding of these exotic hadrons has not yet reached a consensus and is still under intense scientific debates, they deserve further investigation as candidates for tetraquark states.
Recently, there is increasing experimental evidence proving the existence of possible hidden heavy tetraquark states. In 2020, the first candidate for a charged hidden-charm tetraquark state with strangeness, $Z_{cs}(3985)^{-}$, was discovered in the processes of $e^{+}e^{-}\rightarrow K^{+}D_{s}^{-}D^{\ast 0}$ and $K^{+}D_{s}^{\ast-}D^{0}$ by the BESIII collaboration Ablikim et al. (2021). Its non-zero charge characteristic enables it to be clearly distinguished from hybrid or charmonium. Theoretically, $Z_{cs}(3985)^{-}$ is generally interpreted as tetraquark state $cs\bar{c}\bar{u}$ with a quantum number assignment $J^{P}=1^{+}$ Meng et al. (2020); Karliner and Rosner (2021); Giron et al. (2021); Guo et al. (2023); Liu et al. (2025). Later in 2021 the LHCb collaboration reported two hidden-charm exotic structures in the decay $B^{+}\rightarrow J/\psi\phi K^{+}$, referred to as $Z_{cs}(4000)^{+}$ and $Z_{cs}(4220)^{+}$ Aaij et al. (2021). Their quark composition is thought to be $cu\bar{c}\bar{s}$, and their spin-parity quantum quantum numbers are determined to be $J^{P}=1^{+}$ and $1^{+}$ (or $1^{-}$), respectively Aaij et al. (2021). However, theoretical studies suggest that the $Z_{cs}(4000)^{+}$ and $Z_{cs}(4220)^{+}$ may be candidates for charmonia A. and Dhir (2024); Bokade and Bhaghyesh (2025), tetraquarks Li et al. (2024); Vogt (2024); Yu et al. (2024), or molecules Yang et al. (2021); Chen (2024). It is worth noting that the BESIII collaboration recently observed a new $Z_{cs}(4123)^{-}$ in the processes of $e^{+}e^{-}\rightarrow K^{+}D_{s}^{\ast-}D^{\ast 0}+c.c.^{\ast}$, and its quantum number assignment may be $J^{P}=1^{+}$ Ablikim et al. (2023). The discovery of the $Z_{cs}(4123)^{-}$ may add a new member to the family of charged hidden-charm tetraquark with strangeness. In 2023, A near-threshold peaking structure $X(3960)$ is observed in the $D_{s}^{+}D_{s}^{-}$ invariant mass spectrum by the LHCb collaboration, and the quantum number assignment $J^{PC}=1^{++}$ is favored Aaij et al. (2023). Research shows that the nature of $X(3960)$ may also be a tetraquark $cs\bar{c}\bar{s}$ Badalian and Simonov (2023); Hoffer et al. (2024); Chen et al. (2025). Obviously, these exotic hadrons all have the common components $Q\bar{Q}q\bar{q}$, and figuring out their nature will help us better understand QCD.
In this article, we apply the BO approximation to calculate the mass spectra and spatial sizes of tetraquark $Q\bar{Q}q\bar{q}$. The heavy $Q\bar{Q}$ forms a color octet, which combines with a light $q\bar{q}$ octet to produce a color-singlet tetraquark, that is, the color coupling is represented as $8\otimes 8\rightarrow 1$. The light quark (or antiquark) interactions are treated as perturbations to the dominant heavy-quark potential, allowing the Schrödinger equation for the heavy quark separation to be solved numerically. Given the same amount of momentum, the heavy quarks move much more slowly than the light quarks. The calculation of diquark-antidiquark paradigm was extended to tetraquarks. Using $X(3872)$ as input, other possible tetraquark states were studied.
This paper is organized as follows. In Sec. II, We give a brief introduction of the tetraquark Hamiltonian based on the BO approximation. We obtain the BO potential using the perturbation method and construct the basis vector wave functions of the tetraquark system $Q\bar{Q}q\bar{q}$. In Sec. III, we determine the model parameters, such as the masses of quarks and the color coupling strengths. The mass spectra and spatial sizes of various hidden heavy tetraquark states are calculated and discussed in this section. We summarize in Sec. IV. The nature units $c=\hbar=1$ are used throughout.
II The Born-Oppenheimer theoretical model
II.1 Hamiltonian in the Born-Oppenheimer approximation
For the hidden heavy flavor tetraquark system, it is theoretically difficult to achieve a direct analytical solution. We can use BO approximation in non-relativistic quantum mechanics to handle the degrees of freedom of heavy and light quarks separately. The Hamiltonian of tetraquark system reads
$\begin{split}H=&\sum_{i=1}^{4}\left(\frac{p_{i}^{2}}{2m_{i}}\right)+V(x_{A,B})\\ &+V_{I}(x_{A,B},x_{1,2})+V_{\text{conf}},\end{split}$ (1)
where the coordinates and masses of the heavy quarks are denoted as $x_{A,B}$ and $M$, the coordinates and masses of the light quarks are $x_{1,2}$ and $m$. $\sum_{i=1}^{4}\left(\frac{p_{i}^{2}}{2m_{i}}\right)$ represents the kinetic energy of the quarks. $V(x_{A,B})$ describes the interaction between the heavy quarks. $V_{I}(x_{A,B},x_{1,2})$ describes the interactions between the heavy quarks, light quarks, and the light quarks themselves. $V_{\text{conf}}$ describes the color confinement in the singlet state of the hidden-flavor tetraquark system.
In the quark model, the spatial degree of freedom can be encoded in the model parameters. the Hamiltonian for the $S$-wave tetraquark can be written in a simplified form as Maiani et al. (2005, 2019b)
$H=\sum_{i=1}^{4}m_{i}+E+H_{ss},$ (2)
$H_{ss}=\sum_{i<j}2\kappa_{ij}(\mathbf{s}_{i}\cdot\mathbf{s}_{j}),$ (3)
where $m_{i}$ represents the constituent mass of each quark, $E$ is the energy eigenvalue of the tetraquark system obtained based on the BO approximation, and $H_{ss}$ corresponds to the color-spin hyperfine interaction.
II.2 Born-Oppenheimer potention
In the BO approximation, with the positions of the heavy quarks assumed to be fixed, we first consider the Schrödinger equation for the light quarksMaiani et al. (2019b)
$\begin{split}\left(\sum_{\text{light}}\frac{p_{i}^{2}}{2m}\right)f&+V_{I}(x_{A,B},x_{1,2})f\\ &=\mathcal{E}(x_{A,B})f,\end{split}$ (4)
where $f=f(x_{A,B},x_{1,2})$ is the eigenfunction for the light quarks. $\mathcal{E}(x_{A,B})$ is the eigenvalue dependent on the positions of the heavy quarks $x_{A,B}$. $V_{I}(x_{A,B},x_{1,2})$ describes the interaction between the light and heavy quarks. The total wavefunction can be written as
$\Psi=\psi(x_{A,B})f(x_{A,B},x_{1,2}),$ (5)
where $\psi$ describes the heavy quarks, and $f$ describes the light quarks. Under the Born-Oppenheimer approximation, we apply $H$ to $\Psi$ , assuming that the wavefunction $f$ for the light quarks is sensitive to changes in the positions of the heavy quarks, so $\nabla_{i}f$ cannot be neglected. But the wavefunction $\psi$ for the heavy quarks changes slowly with respect to the positions of the heavy quarks, so $\nabla_{i}\psi$ can be approximated as zero.
Thus, for the kinetic term
$-\hbar^{2}\nabla_{i}^{2}\Psi=-\hbar^{2}\left(f\nabla_{i}^{2}\psi+2(\nabla_{i}\psi)\cdot(\nabla_{i}f)+\psi\nabla_{i}^{2}f\right),$ (6)
it can be approximated as
$-\hbar^{2}\nabla_{i}^{2}\Psi\approx-\hbar^{2}\left(f\nabla_{i}^{2}\psi+\psi\nabla_{i}^{2}f\right).$ (7)
We can rewrite the Schrödinger equation as
$\left(\sum_{\text{heavy}}\frac{p_{i}^{2}}{2M}+V_{\text{BO}}(x_{A,B})\right)\psi=E\psi,$ (8)
with
$V_{\text{BO}}(x_{A,B})=V(x_{A,B})+\mathcal{E}(x_{A,B})+V_{\text{conf}}.$ (9)
Next, we will conduct a detailed analysis of the specific form of $V_{\text{BO}}$. We recall that the color coupling between any pair of particles in color representation $R$ is given by
$V_{C}(r)=\lambda_{q_{1}q_{2}}(\mathbf{R})\frac{\alpha_{s}}{r},$ (10)
where $q_{1}$ and $q_{2}$ represent the irreducible representations of the quark pair.
For hidden heavy flavor tetraquarks, the color of the two heavy quarks after coupling can be either $1$ or $8$. In the first case, however, the interaction between $Q\bar{Q}$ and $q\bar{q}$ pairs occurs through the exchange of color singlets. This leads to a model dominated by nuclear-like forces, ultimately resulting in the formation of hadrocharmonium Dubynskiy and Voloshin (2008). Therefore, in our calculation, we exclude the scenario where $Q\bar{Q}$ couples in the color state $1$. So $\lambda_{Q\bar{Q}}$ and $\lambda_{q\bar{q}}$ are given by
$\lambda_{q_{1}q_{2}}(\mathbf{R})=\frac{1}{2}\left[C_{2}(\mathbf{R})-C_{2}(q_{1})-C_{2}(q_{2})\right],$ (11)
where $C_{2}$ is the second Casimir operator. We note the following results:
$C_{2}(1)$ $=0,\quad C_{2}(\mathbf{R})=C_{2}(\mathbf{\bar{R}}),$ (12)
$C_{2}(3)$ $=\frac{4}{3},\quad C_{2}(6)=\frac{10}{3},\quad C_{2}(8)=3.$
In order to derive $\mathcal{E}(x_{A,B})$, we need to solve Schrödinger Eq.(4). Given that the impact of heavy quarks on light quarks is non-negligible, We adopt a method similar to the treatment of hydrogen molecules to handle $Q\bar{Q}q\bar{q}$. Specifically, we model the hidden heavy flavor tetraquarks as two combined hydrogen-like atoms. In this model, a light particle is paired exclusively with one heavy particle, forming two symmetrical orbits. The total energy of the two orbits is $E_{0}$, and the interactions not included in the orbits are treated as perturbations, denoted as $\Delta E$. This gives rise to two possible cases, $Qq$ and $\bar{Q}\bar{q}$, and $\bar{Q}q$ and $Q\bar{q}$. The second case leads to a molecular state of two meson-like combinations, which is not considered in our analysis. We focus only on the tightly bound states formed by the $Qq$ and $\bar{Q}\bar{q}$ combinations.To analyze the interaction between the orbit $Qq$, we introduce the Cornell potential, a prototype for non-relativistic quark interactions.
$V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+\sigma r+V_{0},$ (13)
where $r$ is the effective radius of the hadron, $\alpha_{s}$ is the QCD coupling constant, $\sigma$ is the QCD string tension, and $V_{0}\simeq-0.3\,\text{GeV}$ is a constant.
But the range of the Cornell potential is limited to heavy quarkonium or interactions between quarks with similar masses. For other cases, such as $qQ$, we modify the formula as follows
$V_{q_{1}q_{2}(\bar{q})}(r)=\lambda_{q_{1}q_{2}}\frac{\alpha_{S}}{r}+\frac{3|\lambda_{q_{1}q_{2}}|}{4}\sigma r+V_{0},$ (14)
where $\sigma$ is the string tension extracted from the charmonium spectrum. As for $\lambda_{q_{1}q_{2}}$, in the tetraquark system, $q_{1}$ and $q_{2}$ are in a superposition of two $SU(3)_{c}$ representations with respective amplitudes $a$ and $b$. then we have
$T=a\left[(q_{1}q_{2})_{\mathbf{R}_{1}}\cdots\right]_{\mathbf{R}_{1}}+b\left[(q_{1}q_{2})_{\mathbf{R}_{2}}\cdots\right]_{\mathbf{R}_{2}},$ (15)
$\lambda_{q_{1}q_{2}}=a^{2}\lambda_{q_{1}q_{2}}(\mathbf{R}_{1})+b^{2}\lambda_{q_{1}q_{2}}(\mathbf{R}_{2}).$ (16)
For the $Qq$ orbit, We use the expression from (14) to represent the potential between the $Qq$ quarks,
$V_{Qq}=-\frac{1}{3}\frac{\alpha_{S}}{r}+\frac{1}{4}\sigma+V_{0}.$ (17)
The radial variation wave function is given by
$R(r)=\frac{A^{3/2}}{\sqrt{\pi}}e^{-Ar}.$ (18)
Substituting the model parameters, we can obtain the ground state and the corresponding wave function by using variational calculation.
Due to the high symmetry of the tetraquark system, the wave functions of the two non-interacting orbits can be expressed as
$f(1,2)=\psi(1)\phi(2)=R(|x_{1}-x_{A}|)R(|x_{2}-x_{B}|).$ (19)
The ground state energy corresponding to the wave function is
$E_{0}=2(\langle H\rangle_{\min}+V_{0}).$ (20)
Next, we calculate the perturbation term as
$\begin{split}H_{\text{pert}}=&-\frac{7}{6}\alpha_{S}\left(\frac{1}{|x_{1}-x_{B}|}+\frac{1}{|x_{2}-x_{A}|}\right)\\ &+\frac{1}{6}\alpha_{S}\frac{1}{|x_{1}-x_{2}|}.\end{split}$ (21)
The perturbative energy is given by
$\Delta E=\langle f|H_{\text{pert}}|f\rangle,$ (22)
Further calculation yields
$\Delta E=-\frac{7}{6}\alpha_{S}2I_{1}(r_{AB})+\frac{1}{6}\alpha_{S}I_{4}(r_{AB}),$ (23)
where the functions $I_{1}(r)$ and $I_{4}(r)$ are given by
$I_{1}(r_{AB})=\int d^{3}\xi|\psi(\xi)|^{2}\frac{1}{|\xi-x_{B}|},$ (24)
$I_{4}(r_{AB})=\int d^{3}\xi d^{3}\eta|\psi(\xi)|^{2}|\phi(\eta)|^{2}\frac{1}{|\xi-\eta|}.$ (25)
We obtain the final energy as
$\mathcal{E}(x_{A,B})=E_{0}+\Delta E(x_{A,B}).$ (26)
Finally, we consider the color confinement of the entire tetraquark system. Recall that $3\otimes 3=\bar{3}\oplus 6$, and we observe that the $cq$ orbit is not a singlet, implying it is affected by the confinement potential. We need to include a linear string tension to simulate the confinement effect, which is given by
$V_{\text{conf}}(r)=\sigma_{T}\times(r-R_{0})\times\theta(r-R_{0}).$ (27)
$R_{0}$ is the relative distance between quarks within a hadron. In principle, $R_{0}$ should be treated as a free parameter, and its value can be determined by experimental phenomena related to the tetraquark states. For parameter $\sigma_{T}$, we use the Fierz identity to express the tetraquark state $T=|(c\bar{c})_{\bm{8}}(\bar{q}q)_{\bm{8}}\rangle_{\bm{1}}$ as
$T=\sqrt{\frac{2}{3}}|(Qq)_{\bar{\bm{3}}}(\bar{Q}\bar{q})_{\bm{3}}\rangle_{\bm{1}}-\sqrt{\frac{1}{3}}|(Qq)_{\bm{6}}(\bar{Q}\bar{q})_{\bar{\bm{6}}}\rangle_{\bm{1}}.$ (28)
Due to the gluon screening effect, the diquark in the six-dimensional representation ($6$) adds a component to the three-dimensional representation ($\bar{3}$), making $\sigma_{T}$ closer to $\sigma$. For simplicity, we adopt $\sigma_{T}=\sigma$. Thus, the final potential is
$\begin{split}V_{\text{BO}}(r)=&\frac{1}{6}\alpha_{S}\frac{1}{r_{AB}}+E_{0}+\Delta E(x_{A,B})\\ &+\sigma\times(r-R_{0})\times\theta(r-R_{0}).\end{split}$ (29)
II.3 Wave function
Now we construct the basis vector wave functions of the tetraquark state in spin space. And the total spin of all possible $S$-wave tetraquark states be 0, 1, and 2. For $J^{P(C)}=0^{+(+)}$, there are two spin-zero states,
$|0^{+(+)}\rangle$ $=|0_{Qq},0_{\bar{Q}\bar{q}};J=0\rangle,$ (30)
$|0^{+(+)}\rangle$ $=|1_{Qq},1_{\bar{Q}\bar{q}};J=0\rangle.$
The (+/)(+/-)( + / - ) represents charge conjugation when the tetraquark state is in flavor symmetry, and has the same meaning as below unless otherwise specified.
For $J=1$ with positive parity, there are three states,
$|A\rangle$ $=|0_{Qq},1_{\bar{Q}\bar{q}};J=1\rangle,$ (31)
$|B\rangle$ $=|1_{Qq},0_{\bar{Q}\bar{q}};J=1\rangle,$
$|C\rangle$ $=|1_{Qq},1_{\bar{Q}\bar{q}};J=1\rangle.$
Under charge conjugation, $|A\rangle$ and $|B\rangle$ are exchanged, while $|C\rangle$ has odd parity. Therefore, the $1^{+}$ composite state consists of one $C$-even state and two $C$-odd states,
$|1^{+(+)}\rangle$ $=\frac{1}{\sqrt{2}}(|A\rangle+|B\rangle),$ (32)
$|1^{+(-)}\rangle$ $=\frac{1}{\sqrt{2}}(|A\rangle-|B\rangle),$
$|1^{+(-)^{\prime}}\rangle$ $=|C\rangle.$
We can analyze the specific spin values of these states for both $Q\bar{Q}$ and $q\bar{q}$ quark pairs. Since $J=1$, it is impossible to have a state with zero spin; among the remaining states, only the $C=+$ state has spin 1 for both quarks. Therefore, the state $|1^{+(+)}\rangle$ in Eq.(32) has a definite $Q\bar{Q}$ spin value of $S_{Q\bar{Q}}=1$. For $J^{PC}=2^{+(+)}$,
$|2^{+(+)}\rangle=|1_{Qq},1_{\bar{Q}\bar{q}};J=2\rangle.$ (33)
The $2^{+(+)}$ state also has $S_{Q\bar{Q}}=1$. Therefore, in order to solve the mass spectra and wave functions of the tetraquark states, we only need to substitute the basis vectors into Eq.(2) and diagonalize them.
III Results and discussion
III.1 Model parameters
In numerical calculations, we take the values of $\alpha_{S}$ and $\sigma$, based on the charmonium scale obtained from lattice calculations Kawanai and Sasaki (2012),
$\alpha_{S}(2M_{c})=0.30,\quad\sigma=0.15\,\text{GeV}^{2}.$ (34)
At the $B_{c}$ meson and bottom-charm quark mass scale, we use the same string tension $\sigma$ and $\alpha_{S}$ running via the two-loop beta function, yielding
$\alpha_{S}(M_{c}+M_{b})=0.24,\quad\alpha_{S}(2M_{b})=0.21.$ (35)
Table 1: Constituent quark masses derived from the $S$-wave mesons (first row) and the baryons (second row).
$q$ $s$ $c$ $b$
Mass1 (MeV) 305 490 1670 5005
Mass2 (MeV) 362 546 1721 5044
Table 2: Spin-spin couplings for quark-antiquark pairs in color singlet from the hyperfine splittings of $S$-wave mesons (first row). The values in the second row show the approximate scaling of the couplings with inverse masses (from the meson spectra).
$q\bar{q}$ $s\bar{q}$ $s\bar{s}$ $c\bar{q}$ $c\bar{s}$ $c\bar{c}$
$(\kappa_{ij})_{\bm{0}}$ (MeV) 315 195 121 70 72 59
$(\kappa_{ij})_{\bm{0}}m_{i}m_{j}$ (GeV3) 0.029 0.029 0.036 0.059 0.16
Table 3: Spin-spin couplings for quark-quark pairs in color $\bar{\bm{3}}$ state from $S-$wave baryons. One gluon exchange implies $(\kappa_{ij})_{\bar{\bm{3}}}=1/2(\kappa_{ij})_{\textbf{0}}$. The values in the second row show the approximate scaling of the couplings with inverse masses (from the baryon spectra).
$qq$ $sq$ $cq$ $cs$
$(\kappa_{ij})_{\bar{\bm{3}}}$ (MeV) 103 64 22 25
$(\kappa_{ij})_{\bar{\bm{3}}}m_{i}m_{j}$ (GeV3) 0.014 0.013 0.014 0.024
The numerical solution of the Schrödinger equation will yield the energy eigenvalues $E$ of the tetraquark wave function. However, there still exists a constant $V_{0}$ in Eq. (20) that influences the exact value of the eigenvalue. For tetraquark states, we can approximate the boundary conditions to derive the relationship between $V_{0}$ and $\langle H\rangle_{\text{min}}$, but for hidden heavy flavor tetraquarks states, the boundary conditions are difficult to resolve. We trace back to the original meaning of $V_{0}$, which originates from the Cornell potential as a fitting constant in both experimental and theoretical models. In our analysis, we utilize the $X(3872)$ as an input parameter for the $c\bar{c}q\bar{q}$ system. Through this approach, we determine the optimal solution for $V_{0}=-0.315$ GeV. This value is consistent with the Cornell potential model. We can write the complete expression for the $c\bar{c}q\bar{q}$ mass as
$M(T)=2(M_{c}+M_{q})+E+H_{cm},$ (36)
and plot the relationship between $M(T)$ and $R_{0}$, see Fig.1.
Figure 1: For $J^{PC}=1^{++}$, the ground state energy of the $c\bar{c}q\bar{q}$ system decreases as the system size $R_{0}$ increases.
Besides, we calculated the mass spectra of the tetraquark state at different $R_{0}$, see Table 4. Among them, when $R_{0}=10$ GeV-1, the mass of the tetraquark state $c\bar{c}q\bar{q}$ with quantum number $1^{++}$ is 3.872 GeV.
Table 4: Mass spectra of tetraquark $c\bar{c}q\bar{q}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{++})$ $M(1^{+-})$ $M(1^{++})$ $M(2^{++})$
$R_{0}=6$ 3.792, 3.898 3.785, 3.913 3.903 3.982
$R_{0}=8$ 3.770, 3.876 3.763, 3.891 3.881 3.960
$R_{0}=10$ 3.761, 3.867 3.754, 3.882 3.872 3.951
$R_{0}=12$ 3.757, 3.863 3.750, 3.878 3.868 3.947
Using this method, we can estimate the mass spectra of other tetraquark states.
Figure 2: The mass spectra of the $S$-wave tetraquark states with distinct quantum numbers: (a) $c\bar{c}q\bar{q}$, (b) $c\bar{c}q\bar{w}$, and (c) $c\bar{c}s\bar{s}$. The black dashed lines and red dash-dotted lines represent the corresponding meson-meson thresholds.
Figure 3: The mass spectra of the $S$-wave tetraquark states with distinct quantum numbers: (a) $b\bar{b}q\bar{q}$, (b) $b\bar{b}q\bar{s}$, and (c) $b\bar{b}s\bar{s}$. The black dashed lines and red dash-dotted lines represent all possible meson-meson thresholds.
III.2 Hidden-charm tetraquark states with $c\bar{c}q\bar{q}$, $c\bar{c}q\bar{s}$, and $c\bar{c}s\bar{s}$ configurations
III.2.1 The $c\bar{c}q\bar{q}$ configuration
As long as $Q(q)$ and $\bar{Q}(\bar{q})$ are of the same flavor, the tetraquark state has a well-defined $C$-parity. The tetraquark state can strongly decay into a pair of mesons, and in these processes, the $C$-parity is conserved.
(1) The mass spectrum of $c\bar{c}q\bar{q}$ is shown in FIG. 2(a), with an initial state having $J^{P}=0^{+}$ ($J^{PC}=0^{++}$), which is electrically neutral. Their masses are 3761.1 MeV and 3867.2 MeV, respectively. The lowest state, at 3761.1 MeV, is above the $D\bar{D}$ threshold and can decay via S-wave into $D\bar{D}$ and $\eta_{c}\pi$. The state at 3867.2 MeV is also above the $\eta_{c}\pi$ and $D\bar{D}$ thresholds but slightly below the $J/\psi\rho$ threshold. Considering angular momentum and $C$-parity conservation, this state allows strong decays into $\eta_{c}\pi$, $D\bar{D}$, or $J/\psi\rho$ mesons. This state could be a strong contender for the $X(3915)$ candidate. In this paper, we focus only on $S$-wave decays and neglect higher-order decays, such as $P$-wave and $D$-wave, as these typically contribute less.
(2) There are three $S$-wave tetraquark states with $J^{P}=1^{+}$. The electrically neutral tetraquark states can be classified into two categories based on their $C$-parity. Two $S$-wave tetraquark states have $J^{PC}=1^{+-}$. The lowest state, with a mass of 3754.1 MeV, is slightly below the $\eta_{c}\rho$ threshold. It can easily decay into $\eta_{c}$ and $\rho$, as it has a significant component and a small mass difference. Similarly, the state with a mass of 3882.3 MeV is close to the $D\bar{D}^{*}$ threshold, suggesting that it can easily decay into $D\bar{D}^{*}$. Additionally, it can also decay into $\eta_{c}\rho$ or $J/\psi\pi$. It’s encouraging that these decay channels are considered as potential decay routes for $X(3900)$, which strongly supports the candidacy of this state as $X(3900)$. Another $S$-wave tetraquark state with positive $C$-parity is the one with $J^{PC}=1^{++}$. Its mass is 3872 MeV, which is the result we obtained by fitting the experimental value X(3872). In our model, the most likely decay mode for this particle is $J/\psi\rho$, a result consistent with the findings in Braaten and Kusunoki (2005).
(3) There is an $S$-wave tetraquark state with $J^{P}=2^{+}$ ($J^{PC}=2^{++}$) in the electrically neutral state. Its mass is 3951.1 MeV, which is above the $J/\psi\rho$ threshold, making it possible for the state to decay into $J/\psi\rho$. Additionally, we have calculated the masses of tetraquarks with varying sizes, as depicted in Table 4, and the decay analysis method remains consistent with the aforementioned approach.
III.2.2 The $c\bar{c}q\bar{s}$ configuration
The mass spectrum of $c\bar{c}q\bar{s}$ is shown in FIG. 2(b). For comparison, the corresponding meson-meson thresholds are also presented in the figure. For the tetraquark state of $c\bar{c}q\bar{s}$, the $C$-parity is undetermined, so the quantum numbers of the $S$-wave tetraquark state $c\bar{c}q\bar{s}$ are $J^{P}=0^{+}$, $1^{+}$, and $2^{+}$.
(1) There are two $J^{P}=0^{+}$ states, with masses of 3901.1 MeV and 3983.3 MeV, respectively. The lighter state, at 3901.1 MeV, is above the $\eta_{c}K$ and $D\bar{D}_{s}$ thresholds, and can easily decay into $\eta_{c}K$ and $D\bar{D}_{s}$. The state at 3983.3 MeV not only decays naturally into the mesons $\eta_{c}K$ and $D\bar{D}_{s}$, but is also very close to the $J/\psi K^{*}$ threshold, allowing for the possibility of decaying into the meson pair $J/\psi K^{*}$.
(2) There are three $S$-wave tetraquark states with $J^{P}=1^{+}$. Their masses are 3913.1, 3970, and 3983.2 MeV, respectively. The lightest state, with a mass of 3913.1 MeV, is above the thresholds for $J/\psi K$ and $\eta_{c}K^{*}$, and its decay channels are likely to include $J/\psi K$ or $\eta_{c}K^{*}$. Similarly, the state at 3970 MeV also has decay channels that may include $J/\psi K$ or $\eta_{c}K^{*}$. The state at 3983.2 MeV, which lies just below the $J/\psi K^{*}$ threshold and slightly above the $D\bar{D}_{s}^{*}$ threshold, is most likely to decay into $J/\psi K$, $\eta_{c}K^{*}$, $J/\psi K^{*}$, or $D\bar{D}_{s}^{*}$. It is worth mentioning that the states at 3970 and 3983.2 MeV could potentially be candidates for $Z_{cs}(3985)^{-}$ and $Z_{cs}(4000)^{+}$.
(3) For $J^{P}=2^{+}$, there is an $S$-wave tetraquark state with a mass of 3988.9 MeV. Its mass is slightly below the threshold for $D^{*}\bar{D}_{s}^{*}$, allowing it to naturally decay into the meson pairs $J/\psi K$, $\eta_{c}K^{*}$, $J/\psi K^{*}$, or $D\bar{D}_{s}^{*}$. Additionally, we have calculated the masses of tetraquarks with varying sizes, as depicted in TABLE 5, and the decay analysis method remains consistent with the aforementioned approach.
Table 5: Mass spectra of tetraquark $c\bar{c}q\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{+})$ $M(1^{+})$ $M(2^{+})$
$R_{0}=6$ 3.929, 4.011 3.942, 3.998, 4.011 4.017
$R_{0}=8$ 3.909, 3.991 3.922, 3.978, 3.991 3.997
$R_{0}=10$ 3.901, 3.983 3.913, 3.970, 3.983 3.989
$R_{0}=12$ 3.898, 3.980 3.910, 3.967, 3.979 3.986
III.2.3 The $c\bar{c}s\bar{s}$ configuration
The tetraquark configuration $c\bar{c}s\bar{s}$ has a well-defined $C$-parity. The possible quantum numbers are $J^{PC}=0^{++},1^{+-},2^{++}$. The mass spectrum of the tetraquark states with the $c\bar{s}c\bar{s}$ configuration is shown in FIG. 2(c), along with the corresponding meson-meson thresholds.
(1) There are two $S$-wave tetraquark states with $J^{PC}=0^{++}$, having masses of 4032.2 MeV and 4122.3 MeV, respectively. The lowest mass state, at 4032.2 MeV, is above the $D_{s}\bar{D}_{s}$ meson-meson threshold, so its possible decay mode is into $D_{s}\bar{D}_{s}$. The state at 4122.3 MeV, which lies just below the $D_{s}^{*}\bar{D}_{s}^{*}$ threshold, can naturally decay into $D_{s}\bar{D}_{s}$ and $J/\psi\phi$.
(2) There are two $S$-wave tetraquark states with $J^{PC}=1^{+-}$, having masses of 4069.1 MeV and 4102.8 MeV, respectively. The lighter state, at 4069.1 MeV, is well above the $\eta_{c}\phi$ threshold, allowing it to naturally decay into $\eta_{c}\phi$. The state at 4102.8 MeV, in addition to the decay mode into $\eta_{c}\phi$, also has a possible decay mode into $D_{s}\bar{D}_{s}^{*}$. Furthermore, there is an $S$-wave tetraquark state with $J^{PC}=1^{++}$ and a mass of 4088.2 MeV. For this state, both $\eta_{c}\phi$ and $D_{s}\bar{D}_{s}^{*}$ decay modes are allowed. Although several promising exotic $c\bar{c}s\bar{s}$ tetraquark candidates had been identified prior to this, our experimental theoretical results more strongly support $X(4140)$ as a $c\bar{c}s\bar{s}$ tetraquark state.
(3) For $J^{PC}=2^{++}$, we observe an $S$-wave tetraquark state with a mass of 4174.3 MeV. Considering $J^{PC}$ conservation, this state can only decay into the meson pair $J/\psi\phi$. Additionally, we have calculated the masses of tetraquarks with varying sizes, as depicted in Table 6, and the decay analysis method remains consistent with the aforementioned approach.
Table 6: Mass spectra of tetraquark $c\bar{c}s\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{++})$ $M(1^{+-})$ $M(1^{++})$ $M(2^{++})$
$R_{0}=6$ 4.060, 4.150 4.096, 4.130 4.115 4.201
$R_{0}=8$ 4.041, 4.130 4.077, 4.111 4.096 4.182
$R_{0}=10$ 4.032, 4.122 4.069, 4.103 4.088 4.174
$R_{0}=12$ 4.030, 4.120 4.066, 4.101 4.086 4.172
III.3 Hidden-bottom tetraquark states with $b\bar{b}q\bar{q}$, $b\bar{b}q\bar{s}$, and $b\bar{b}s\bar{s}$ configurations
The hidden bottom quark tetraquark states can be realized by converting the charm quarks of the hidden charm tetraquark states into bottom quarks. These hidden bottom quark tetraquark states can be classified into three configurations: $b\bar{b}q\bar{q}$, $b\bar{b}q\bar{s}$, and $b\bar{b}s\bar{s}$. The mass spectrum at size $R_{0}=10GeV^{-1}$ is presented in FIG. 3. The analytical approach is similar to that used for the hidden charm tetraquark states. Therefore, we will not discuss each state in detail, but instead focus on the two states observed experimentally: $Z_{b}(10610)$ and $Z_{b}(10650)$. These states have been observed in the decay channels $\Upsilon(nS)\pi$ and $h_{b}(mP)$, where $n=1,2,3$ and $m=1,2$. Their quantum numbers are likely $J^{P}=1^{+}$, but there is no information regarding their $C$-parity. In our calculations, we prefer these two states to be the tetraquark states $b\bar{b}u\bar{d}$ or $b\bar{b}u\bar{d}$ with quantum number assignment $1^{+-}$, as shown in FIG. 3(a). In addition, we have calculated the masses of hidden-bottom tetraquark states at different sizes, as illustrated in Table 7-9, and the analysis method remains consistent.
Table 7: Mass spectra of tetraquark $b\bar{b}q\bar{q}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{++})$ $M(1^{+-})$ $M(1^{++})$ $M(2^{++})$
$R_{0}=6$ 10.603, 10.458 10.598, 10.455 10.605 10.621
$R_{0}=8$ 10.596, 10.451 10.591, 10.449 10.598 10.614
$R_{0}=10$ 10.593, 10.448 10.588, 10.446 10.595 10.611
$R_{0}=12$ 10.592, 10.447 10.587, 10.445 10.584 10.610
Table 8: Mass spectra of tetraquark $b\bar{b}q\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{+})$ $M(1^{+})$ $M(2^{+})$
$R_{0}=6$ 10.604, 10.694 10.604, 10.681, 10.720 10.684
$R_{0}=8$ 10.597, 10.687 10.598, 10.675, 10.713 10.678
$R_{0}=10$ 10.596, 10.685 10.596, 10.673, 10.712 10.676
$R_{0}=12$ 10.595, 10.685 10.595, 10.672, 10.711 10.676
Table 9: Mass spectra of tetraquark $b\bar{b}s\bar{s}$ states at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{++})$ $M(1^{+-})$ $M(1^{++})$ $M(2^{++})$
$R_{0}=6$ 10.696, 10.830 10.756, 10.834 10.789 10.877
$R_{0}=8$ 10.690, 10.824 10.750, 10.828 10.783 10.871
$R_{0}=10$ 10.688, 10.822 10.748, 10.826 10.781 10.869
$R_{0}=12$ 10.687, 10.821 10.747, 10.825 10.780 10.868
IV summary
In this paper, we investigate the hidden heavy flavor tetraquark states $Q\bar{Q}q\bar{q}$ composed of two heavy quarks and two light quarks. A remarkable feature of these tetraquark states is that they conform to a hydrogen-bond-like structure in quantum chromodynamics vertion, which can be treated using the BO approximation. Subsequently, we systematically calculated the mass spectra of $S$-wave tetraquark states $Q\bar{Q}q\bar{q}$ with various quantum numbers ($J^{P}=0^{+(+)}$, $1^{+(\pm)}$, and $2^{+(+)}$), and analyzed their decay channels.
Based on the fact that heavy quarks are much heavier than light quarks, we assume that the wave functions of heavy and light quarks in the tetraquark state $Q\bar{Q}q\bar{q}$ can be treated separately. We consider the interaction between light quarks as a perturbation and introduce the BO potential. In specific calculations, the BO potential can be divided into three parts: the QCD Coulomb interaction, the eigenenergy of the light quark systems, and a linear term introduced to account for the color confinement of the tetraquark state. Analogous to the treatment of $H_{2}$ in QED, we solve the two-body Schrödinger equation to obtain the zeroth-order approximation of the energy occupied by light quarks within the tetraquark state, and thus obtain the eigenvalues of the light quark Hamiltonian. For better comparison, we use X(3872) as input to determine $R_{0}=10\,\text{GeV}^{-1}$, and study other hidden heavy flavor tetraquark states.
Our calculation results indicate that the exotic hadrons $X(4140)$, $Y(3940)$, $Z_{b}(10610)$, and $Z_{b}(10650)$ remain strong candidates for tetraquark states. Besides, we propose that $Z_{cs}(3985)^{-}$, $Z_{cs}(4000)^{+}$, and $Z_{cs}(4123)^{-}$ may be partner states of the tetraquark state $cu\bar{c}\bar{s}/cs\bar{c}\bar{u}$ with quantum number $J^{P}=1^{+}$. The $X(3960)$ may be a tetraquark state $cs\bar{c}\bar{s}$ with larger spatial sizes and quantum numbers $J^{PC}=1^{++}$. In addition, we predict several other possible hidden heavy flavor tetraquark states, and look forward to further experimental and theoretical research. Finally, it is worth emphasizing that although the difference in quark masses may have some impact on the mass spectra of tetraquark states, this hydrogen-bond-like structure in quantum chromodynamics should not be overlooked.
Acknowledgements.
The authors would like to thank Yanmei Xiao and Lei Liu for helpful comments and discussions. This work was supported by the Scientific Research Foundation of Chengdu University of Technology under grant No.10912-KYQD2022-09557.